Determine Ka of the weak acid HX knowing that 0.10M solution of LiX has pH of 8.90

Post-Laboratory Questions and Exercises DUE AFTER COMPLETING LAB. ANSWER IN SPACE PROVIDED. 1. Determine the K, of the weak acid HX knowing that a 0.10 M solution of LiX has pH of 8.90.
Answer #1

LiX is salt of strong base and weak acid.

for such salts

pH = 1/2 [pKw + pKa + log C]

8.90 = 1/2 [14 + pKa + log 0.10]

17.80 = 13 + pKa

pKa = 4.80

Ka = 10-pKa = 10-4.8

Ka = 1.60 x 10-5